OpenStudy (nadiahossain):

2.4*10^6 was rewritten as 24*10^5. Explain why those expressions are equivalent.

OpenStudy (anonymous):

this tutorial will help you http://openstudy.com/study#/updates/56856b62e4b032ed60def0df

OpenStudy (dayakar):

2.4 *10^6= 24/10 * 10^6 can u simplify it get the the answer

OpenStudy (anonymous):

\[2.4*10^6=\frac{ 24 }{ 10 }*10,00,000=24,00,000\] \[24*10^5=24*1,00,000=24,00,000\]